# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ 5 Conclusion
Since ${\cal X}$ is compact under Assumption (A1), it can be verified that $\boldsymbol{\Xi}({\cal X})$ is also compact under the topology of weak convergence, that is, $\boldsymbol{\xi}_{n}$ converges weakly to $\boldsymbol{\xi}_{0}$ , as $n$ goes to $\infty$, if and only if $\lim_{n\rightarrow\infty}\int_{\cal X}f({\mathbf{x}})\boldsymbol{\xi}_{n}(d{\mathbf{x}})=\int_{\cal X}f({\mathbf{x}})\boldsymbol{\xi}_{0}(d{\mathbf{x}})$ for all bounded continuous function $f$ on ${\cal X}$ (Billingsley, 1999; Fedorov and Leonov, 2014). Further with Assumption (A2), $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$ due to Lemma S2.1, and must be bounded due to the compactness of ${\cal X}$, then ${\mathbf{F}}_{\rm SEW}$ as a map from $\boldsymbol{\Xi}({\cal X})$ to $\mathbb{R}^{p\times p}$ is also bounded continuous, and thus its image ${\cal F}_{\rm SEW}({\cal X})$ is also compact. Similarly with Assumption (A3) and Lemma S2.2, ${\mathbf{F}}_{\rm EW}$ is a bounded continuous map from $\boldsymbol{\Xi}({\cal X})$ to $\mathbb{R}^{p\times p}$, and ${\cal F}_{\rm EW}({\cal X})$ is compact as well.
As a summary, we have the following lemma:
Lemma S2.3.
Under Assumptions (A1) and (A2), ${\cal F}_{\rm SEW}({\cal X})$ is convex and compact; under Assumptions (A1) and (A3), ${\cal F}_{\rm EW}({\cal X})$ is convex and compact as well.
Theorem S2.1.
Proof of Theorem S2.1.
Under Assumptions (A1) and (A2), following the proof of the lemma in Section 1.26 of Pukelsheim (1993), it can be verified that $\{{\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}\}\subset\mathbb{R}^{p\times p}$, as the convex hull of the set $\{\hat{E}\{{\mathbf{F}}(\mathbf{x},\boldsymbol{\Theta})\}\mid\mathbf{x}\in{\cal X}\}$ with dimension $p(p+1)/2$ due to their symmetry, is convex and compact. Due to Lemma S2.3 and Carathéodory’s Theorem (see, for example, Theorem 2.1.1 in Fedorov (1972)), it can be further verified that for any matrix ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ with $\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})$, there exists a $\boldsymbol{\xi}_{0}\in\boldsymbol{\Xi}$ with at most $p(p+1)/2+1$ design points, such that, ${\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi}_{0})={\mathbf{F}}$. When ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ is a boundary point (that is, any open set of ${\cal F}_{\rm SEW}({\cal X})$ containing ${\mathbf{F}}$ contains both matrices inside and outside ${\cal F}_{\rm SEW}({\cal X})$), the number of design points in $\boldsymbol{\xi}_{0}$ can be at most $p(p+1)/2$.
To characterize D-optimality for EW designs, we define the objective function
$\Psi({\mathbf{F}})=-\log|{\mathbf{F}}|$
for ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ or ${\cal F}_{\rm EW}({\cal X})$. Then $f_{\rm SEW}(\boldsymbol{\xi})=\exp\{-\Psi(\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\})\}$, and $f_{\rm EW}(\boldsymbol{\xi})=\exp\{-\Psi(E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\})\}$. Note that minimizing $\Psi({\mathbf{F}})$ for ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ or ${\cal F}_{\rm EW}({\cal X})$ is equivalent to maximizing $f_{\rm SEW}(\boldsymbol{\xi})$ or $f_{\rm EW}(\boldsymbol{\xi})$ for $\boldsymbol{\xi}\in\boldsymbol{\Xi}$, respectively, due to Theorem S2.1.
@@
According to Section 2.4.2 in Fedorov and Leonov (2014), for D-optimality, $\Psi$ always satisfies their Assumptions (B1), (B2), and (B4). In our notations, we let $\boldsymbol{\Xi}(q)$ denote $\{\boldsymbol{\xi}\in\boldsymbol{\Xi}\mid|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|\geq q\}$ for sample-based EW D-optimality, or $\{\boldsymbol{\xi}\in\boldsymbol{\Xi}\mid|E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|\geq q\}$ for integral-based EW D-optimality. We also denote ${\mathbf{F}}(\boldsymbol{\xi})=\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for sample-based EW D-optimality, and ${\mathbf{F}}(\boldsymbol{\xi})=E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for integral-based EW D-optimality. Similarly, ${\mathbf{F}}_{\mathbf{x}}=\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ for sample-based EW D-optimality, and ${\mathbf{F}}_{\mathbf{x}}=E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ for integral-based EW D-optimality. Then the corresponding assumptions are provided as below.
@@
S3 First-order Derivative of Sensitivity Function
To implement the EW ForLion algorithm (Algorithm 1) for EW D-optimal designs, if there is at least one continuous factor, we need to calculate the first-order derivative of the sensitivity function $d({\mathbf{x}},\boldsymbol{\xi})$ (Huang et al., 2024).
Following the simplified notation in Section 3.2, we denote
${\mathbf{F}}(\boldsymbol{\xi})$ $=$ $\left\{\begin{array}[]{cl}E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}&\mbox{, under integral-based EW D-optimality;}\\ \hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}&\mbox{, under sample-based EW D-optimality,}\end{array}\right.$
${\mathbf{F}}_{\mathbf{x}}$ $=$ $\left\{\begin{array}[]{cl}E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}&\mbox{, under integral-based EW D-optimality;}\\ \hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}&\mbox{, under sample-based EW D-optimality,}\end{array}\right.$
Then $d({\mathbf{x}},\boldsymbol{\xi})$ in Theorem 2.1 can be simplified to $d({\mathbf{x}},\boldsymbol{\xi})={\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})$.
Similar to Theorem 3 in Huang et al. (2024), we obtain the following theorem for multinomial logistic models (MLM) under EW D-optimality.
Theorem S3.1.
For an MLM model under Assumptions (A1) and (B3), consider $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ satisfying $|{\mathbf{F}}(\boldsymbol{\xi})|>0$. If we denote the $p\times p$ matrix
${\mathbf{F}}(\boldsymbol{\xi})^{-1}=\left[\begin{array}[]{ccc}\mathbf{E}_{11}&\cdots&\mathbf{E}_{1J}\\ \vdots&\ddots&\vdots\\ \mathbf{E}_{J1}&\cdots&\mathbf{E}_{JJ}\end{array}\right]$
with submatrix $\mathbf{E}_{st}\in\mathbf{R}^{p_{s}\times p_{t}}$ (we denote $p_{J}=p_{c}$ for simplicity), then
$d(\mathbf{x},\boldsymbol{\xi})$ $=$ ${\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})$
$=$ $\sum_{t=1}^{J-1}u_{tt}^{\mathbf{x}}\left(\mathbf{h}_{t}^{\mathbf{x}}\right)^{T}\mathbf{E}_{tt}\mathbf{h}_{t}^{\mathbf{x}}+\sum_{s=1}^{J-1}\sum_{t=1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{c}^{\mathbf{x}}\right)^{T}\mathbf{E}_{JJ}\mathbf{h}_{c}^{\mathbf{x}}$
$+$ $2\sum_{s=1}^{J-2}\sum_{t=s+1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{t}^{\mathbf{x}}\right)^{T}\mathbf{E}_{st}\mathbf{h}_{s}^{\mathbf{x}}+2\sum_{s=1}^{J-1}\sum_{t=1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{c}^{\mathbf{x}}\right)^{T}\mathbf{E}_{sJ}\mathbf{h}_{s}^{\mathbf{x}}\ ,$
with ${\mathbf{h}}_{j}^{\mathbf{x}}={\mathbf{h}}_{j}({\mathbf{x}})$, ${\mathbf{h}}_{c}^{\mathbf{x}}={\mathbf{h}}_{c}({\mathbf{x}})$, and
$u_{st}^{\mathbf{x}}$ $=$ $\left\{\begin{array}[]{cl}E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}&\mbox{, under integral-based EW D-optimality;}\\ \hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}&\mbox{, under sample-basedEW D-optimality,}\end{array}\right.$
as in Example 3.1 for simplicity.
Example S3.1.
Multinomial logistic models (MLM):To calculate the first-order derivative of $d({\mathbf{x}},\boldsymbol{\xi})$ under MLMs, we follow Appendix C in Huang et al. (2024) after
  • (i)
    ${\mathbf{F}}(\boldsymbol{\xi})$ is replaced with $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for each $\boldsymbol{\xi}\in\boldsymbol{\Xi}$;
  • (ii)
    ${\mathbf{F}}_{\mathbf{x}}$ is replaced with $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ for each ${\mathbf{x}}\in{\cal X}$;
  • (iii)
    ${\mathbf{U}}_{\mathbf{x}}=(u_{st}^{\mathbf{x}})_{s,t=1,\ldots,J}$ is replaced by $\left(E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}\right)_{s,t=1,\ldots,J}$ or $\left(\hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}\right)_{s,t=1,\ldots,J}$ (see also Example 3.1);
  • (iv)
    If $k\geq 1$, for $i=1,\ldots,k$, $\frac{\partial{\mathbf{U}}_{\mathbf{x}}}{\partial x_{i}}=\left(\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}\right)_{s,t=1,\ldots,J}$ is replaced by
    $\left(\frac{\partial E\left\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\right\}}{\partial x_{i}}\right)_{s,t=1,\ldots,J}=\left(\frac{\partial}{\partial x_{i}}\int_{\boldsymbol{\Theta}}u_{st}^{\mathbf{x}}(\boldsymbol{\theta})Q(d\boldsymbol{\theta})\right)_{s,t=1,\ldots,J}$
    or
    $\left(\hat{E}\left\{\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\boldsymbol{\Theta})\right\}\right)_{s,t=1,\ldots,J}=\left(\frac{1}{B}\sum_{j=1}^{B}\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\hat{\boldsymbol{\theta}}_{j})\right)_{s,t=1,\ldots,J}\ .$ (S.7)
Note that $\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\hat{\boldsymbol{\theta}}_{j})$ in (S.7) denotes $\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}$ at $\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}_{j}$ One can use the formulae in Appendix C of Huang et al. (2024) to calculate it for each $j=1,\ldots,B$. $\Box$
Example S3.2.
Generalized linear models (GLM):Under GLMs, the sensitivity function
$d({\mathbf{x}},\boldsymbol{\xi})=E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}\left[E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}\right]^{-1}{\mathbf{h}}({\mathbf{x}})$ (S.8)
for $f_{\rm EW}(\boldsymbol{\xi})$ or
$d({\mathbf{x}},\boldsymbol{\xi})=\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}\left[\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}\right]^{-1}{\mathbf{h}}({\mathbf{x}})$ (S.9)
for $f_{\rm SEW}(\boldsymbol{\xi})$. Following the notation in Section S.5 of the Supplementary Material of Huang et al. (2024), we denote ${\mathbf{A}}_{E}=[{\mathbf{X}}_{{\boldsymbol{\xi}}}^{T}E\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{{\boldsymbol{\xi}}}]^{-1}$ and ${\mathbf{A}}_{\hat{E}}=[{\mathbf{X}}_{{\boldsymbol{\xi}}}^{T}\hat{E}\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{{\boldsymbol{\xi}}}]^{-1}$. The first-order derivative of $d({\mathbf{x}},\boldsymbol{\xi})$ with respect to continuous factors ${\mathbf{x}}_{(1)}=(x_{1},\ldots,x_{k})^{T}$ is
$\frac{\partial d({\mathbf{x}},\boldsymbol{\xi})}{\partial{\mathbf{x}}_{(1)}}={\mathbf{h}}({\mathbf{x}})^{T}{\mathbf{A}}_{E}{\mathbf{h}}({\mathbf{x}})\cdot\frac{\partial E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]}{\partial{\mathbf{x}}_{(1)}}+2\cdot E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]\cdot\left\{\frac{\partial{\mathbf{h}}({\mathbf{x}})}{\partial{\mathbf{x}}_{(1)}^{T}}\right\}^{T}{\mathbf{A}}_{E}{\mathbf{h}}({\mathbf{x}})$
for $f_{\rm EW}(\boldsymbol{\xi})$, or
$\frac{\partial d({\mathbf{x}},\boldsymbol{\xi})}{\partial{\mathbf{x}}_{(1)}}$ $=$ ${\mathbf{h}}({\mathbf{x}})^{T}{\mathbf{A}}_{\hat{E}}{\mathbf{h}}({\mathbf{x}})\cdot\frac{\partial\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]}{\partial{\mathbf{x}}_{(1)}}+2\cdot\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]\cdot\left\{\frac{\partial{\mathbf{h}}({\mathbf{x}})}{\partial{\mathbf{x}}_{(1)}^{T}}\right\}^{T}{\mathbf{A}}_{\hat{E}}{\mathbf{h}}({\mathbf{x}})$
$=$ ${\mathbf{h}}({\mathbf{x}})^{T}{\mathbf{A}}_{\hat{E}}{\mathbf{h}}({\mathbf{x}})\cdot\left\{\frac{\partial{\mathbf{h}}({\mathbf{x}})}{\partial{\mathbf{x}}_{(1)}^{T}}\right\}^{T}\cdot\frac{1}{B}\sum_{j=1}^{B}\nu^{\prime}\{{\mathbf{h}}({\mathbf{x}})^{T}{\hat{\boldsymbol{\theta}}}_{j}\}\hat{\boldsymbol{\theta}}_{j}$
for $f_{\rm SEW}(\boldsymbol{\xi})$. $\Box$
Theorem S3.2.
For a GLM model under Assumptions (A1) and (B3), consider $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ satisfying $|\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}|>0$, where ${\mathbf{W}}_{{\boldsymbol{\xi}}}$ is either $E\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}$ for integral-based EW D-optimality or $\hat{E}\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}$ for sample-based EW D-optimality (see Example 3.2). Then $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}[\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}]^{-1}\cdot$ ${\mathbf{h}}({\mathbf{x}})\leq p$ for integral-based D-optimality or $\max_{\mathbf{x}\in\mathcal{X}}\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}[\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}]^{-1}{\mathbf{h}}({\mathbf{x}})\leq p$ for sample-based EW D-optimality.
S4 Proofs of Main Theorems
Proof of Theorem 2.1.
@@
For sample-based EW D-optimality, under Assumptions (A1), (A2), and (B3), due to Lemma S2.3, ${\cal F}_{\rm SEW}({\cal X})$ is convex and compact. Due to Assumption (B3), there exists a sample-based EW D-optimal design $\boldsymbol{\xi}_{*}\in\boldsymbol{\Xi}({\cal X})$. Due to the monotonicity of $\Psi$ (see Assumption (B2)), $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}_{*},\boldsymbol{\Theta})\}$ must be a boundary point of ${\cal F}_{\rm SEW}({\cal X})$. According to Theorem S2.1, there exists a sample-based EW D-optimal design with no more than $p(p+1)/2$ support points.
@@
According to Section 2.4.2 in Fedorov and Leonov (2014), Assumptions (B1), (B2), and (B4) are always satisfied for D-optimality. The rest parts of Theorem 2.1 are direct conclusions of Theorem 2.2 in Fedorov and Leonov (2014). ∎
Proof of Theorem 3.1.
@@
Suppose all the predictor functions ${\mathbf{h}}_{1},\ldots,{\mathbf{h}}_{J-1}$ and ${\mathbf{h}}_{c}$ defined in (5) are continuous with respect to all continuous factors of $\mathbf{x}\in{\cal X}$. Then there exists an $M_{x}>0$ due to the compactness of ${\cal X}$, such that, $\|{\mathbf{h}}_{j}({\mathbf{x}})\|\leq M_{x}$ and $\|{\mathbf{h}}_{c}({\mathbf{x}})\|\leq M_{x}$ , for all $j=1,\ldots,J-1$ and ${\mathbf{x}}\in{\cal X}$.
@@
According to Appendix A of Huang et al. (2024), if we further have a bounded $\boldsymbol{\Theta}$, then there exists an $M_{\eta}>0$, such that, $\|\boldsymbol{\eta}_{\mathbf{x}}\|=\|{\mathbf{X}}_{\mathbf{x}}\boldsymbol{\theta}\|\leq M_{\eta}$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ and ${\mathbf{x}}\in{\cal X}$. Then there exist $10<\delta_{x}<\Delta_{x}<$ 1, such that, $\delta_{x}\leq\pi_{j}^{\mathbf{x}}(\boldsymbol{\theta})\leq\Delta_{x}$ for all $j=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. Then there exists an $M_{u}>0$, such that, $|u_{st}^{\mathbf{x}}(\boldsymbol{\theta})|\leq M_{u}$ for all $s,t=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. According to Theorem 2 in Huang et al. (2024), there exists an $M_{F}>0$, such that, $\|{\mathbf{F}}_{st}^{\mathbf{x}}(\boldsymbol{\theta})\|=\|{\mathbf{F}}_{st}({\mathbf{x}},\boldsymbol{\theta})\|\leq M_{F}$ for all $s,t=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. As a direct conclusion, the Fisher information of an MLM satisfies Assumption (A3). The rest statements are direct conclusions of Theorem 2.1 and Corollary 3.1. ∎
Proof of Theorem 3.2.
@@
Suppose all the predictor functions ${\mathbf{h}}=(h_{1},\ldots,h_{p})^{T}$ are all continuous with respect to $\mathbf{x}\in{\cal X}$, and $\boldsymbol{\Theta}$ is bounded. Then there exist $M_{x}>0$ and $M_{\eta}>0$, such that, $\|{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}\|\leq M_{x}$ for all ${\mathbf{x}}\in{\cal X}$, and $|{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}|\leq M_{\eta}$ for all ${\mathbf{x}}\in{\cal X}$ and $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Since $\nu$ is continuous for commonly used GLMs (see, for example, Table 5 in the Supplementary Material of Huang et al. (2025)), there is an $M_{F}>0$, such that, $|{\mathbf{F}}_{st}({\mathbf{x}},\boldsymbol{\theta})|\leq M_{F}$ for all $s,t=1,\ldots,p$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. As a direct conclusion, the Fisher information of a commonly used GLM satisfies Assumption (A3). The rest statements are direct conclusions of Theorem 2.1 and Corollary 3.1. ∎
Proof of Theorem S3.1.
@@
According to Theorem 2.1, design $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})$ $\leq p$. By applying Theorem 2 and its proof (see their Lemma 1) of Huang et al. (2024), we obtain

### v2

@@ 5 Conclusion
Since ${\cal X}$ is compact under Assumption (A1), it can be verified that $\boldsymbol{\Xi}({\cal X})$ is also compact under the topology of weak convergence, that is, $\boldsymbol{\xi}_{n}$ converges weakly to $\boldsymbol{\xi}_{0}$ , as $n$ goes to $\infty$, if and only if $\lim_{n\rightarrow\infty}\int_{\cal X}f({\mathbf{x}})\boldsymbol{\xi}_{n}(d{\mathbf{x}})=\int_{\cal X}f({\mathbf{x}})\boldsymbol{\xi}_{0}(d{\mathbf{x}})$ for all bounded continuous function $f$ on ${\cal X}$ (Billingsley, 1999; Fedorov and Leonov, 2014). Further with Assumption (A2), $\hat{E}\{{\mathbf{F}}({\mathbf{x}},$ $\boldsymbol{\Theta})\}$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$ due to Lemma 5, and must be bounded due to the compactness of ${\cal X}$, then ${\mathbf{F}}_{\rm SEW}$ as a map from $\boldsymbol{\Xi}({\cal X})$ to $\mathbb{R}^{p\times p}$ is also bounded continuous, and thus its image ${\cal F}_{\rm SEW}({\cal X})$ is also compact. Similarly with Assumption (A3) and Lemma 6, ${\mathbf{F}}_{\rm EW}$ is a bounded continuous map from $\boldsymbol{\Xi}({\cal X})$ to $\mathbb{R}^{p\times p}$, and ${\cal F}_{\rm EW}({\cal X})$ is compact as well.
As a summary of the above arguments, we obtain Lemma 1.
@@
Recall that we denote ${\mathbf{F}}(\boldsymbol{\xi})=\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for sample-based EW D-optimality, and $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for integral-based EW D-optimality; ${\mathbf{F}}_{\mathbf{x}}=\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ for sample-based EW D-optimality, and $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ for integral-based EW D-optimality. Then the same set of assumptions listed in Section 2.4.2 in Fedorov and Leonov (2014) are provided in our notations as below.
@@
According to Section 2.4.2 in Fedorov and Leonov (2014), for D-optimality, Assumptions (B1), (B2), and (B4) are always satisfied. Further discussions have been provided in Section 2.2.
S3 Proofs of Main Theorems
Proof of Theorem 1.
Under Assumptions (A1) and (A2), following the proof of the lemma in Section 1.26 of Pukelsheim (1993), it can be verified that $\{{\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}\}\subset\mathbb{R}^{p\times p}$, as the convex hull of the set $\{\hat{E}\{{\mathbf{F}}(\mathbf{x},\boldsymbol{\Theta})\}\mid\mathbf{x}\in{\cal X}\}$ with dimension $p(p+1)/2$ due to their symmetry, is convex and compact. Due to Lemma 1 and Carathéodory’s Theorem (see, for example, Theorem 2.1.1 in Fedorov (1972)), it can be further verified that for any matrix ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ with $\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})$, there exists a $\boldsymbol{\xi}_{0}\in\boldsymbol{\Xi}$ with at most $p(p+1)/2+1$ design points, such that, ${\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi}_{0})={\mathbf{F}}$. When ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ is a boundary point (that is, any open set of ${\cal F}_{\rm SEW}({\cal X})$ containing ${\mathbf{F}}$ contains both matrices inside and outside ${\cal F}_{\rm SEW}({\cal X})$), the number of design points in $\boldsymbol{\xi}_{0}$ can be at most $p(p+1)/2$.
Proof of Theorem 2.
@@
For sample-based EW D-optimality, under Assumptions (A1), (A2), and (B3), due to Lemma 1, ${\cal F}_{\rm SEW}({\cal X})$ is convex and compact. Due to Assumption (B3), there exists a sample-based EW D-optimal design $\boldsymbol{\xi}_{*}\in\boldsymbol{\Xi}({\cal X})$. Due to the monotonicity of $\Psi$ (see Assumption (B2) in Section S2), $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi}_{*},\boldsymbol{\Theta})\}$ must be a boundary point of ${\cal F}_{\rm SEW}({\cal X})$. According to Theorem 1, there exists a sample-based EW D-optimal design with no more than $p(p+1)/2$ support points.
@@
According to Section 2.4.2 in Fedorov and Leonov (2014), Assumptions (B1), (B2), and (B4) are always satisfied for D-optimality. The rest parts of Theorem 2 are direct conclusions of Theorem 2.2 in Fedorov and Leonov (2014). ∎
Proof of Theorem 3.
@@
Suppose all the predictor functions ${\mathbf{h}}_{1},\ldots,{\mathbf{h}}_{J-1}$ and ${\mathbf{h}}_{c}$ defined in (3.7) are continuous with respect to all continuous factors of $\mathbf{x}\in{\cal X}$. Then there exists an $M_{x}>0$ due to the compactness of ${\cal X}$, such that, $\|{\mathbf{h}}_{j}({\mathbf{x}})\|\leq M_{x}$ and $\|{\mathbf{h}}_{c}({\mathbf{x}})\|\leq M_{x}$ , for all $j=1,\ldots,J-1$ and ${\mathbf{x}}\in{\cal X}$.
@@
According to Appendix A of Huang et al. (2024), if we further have a bounded $\boldsymbol{\Theta}$, then there exists an $M_{\eta}>0$, such that, $\|\boldsymbol{\eta}_{\mathbf{x}}\|=\|{\mathbf{X}}_{\mathbf{x}}\boldsymbol{\theta}\|\leq M_{\eta}$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ and ${\mathbf{x}}\in{\cal X}$. Then there exist $10<\delta_{x}<\Delta_{x}<$ 1, such that, $\delta_{x}\leq\pi_{j}^{\mathbf{x}}(\boldsymbol{\theta})\leq\Delta_{x}$ for all $j=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. Then there exists an $M_{u}>0$, such that, $|u_{st}^{\mathbf{x}}(\boldsymbol{\theta})|\leq M_{u}$ for all $s,t=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. According to Theorem 2 in Huang et al. (2024), there exists an $M_{F}>0$, such that, $\|{\mathbf{F}}_{st}^{\mathbf{x}}(\boldsymbol{\theta})\|=\|{\mathbf{F}}_{st}({\mathbf{x}},\boldsymbol{\theta})\|\leq M_{F}$ for all $s,t=1,\ldots,J$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. As a direct conclusion, the Fisher information of an MLM satisfies Assumption (A3). The rest statements are direct conclusions of Theorem 2 and Corollary 1. ∎
Proof of Theorem 4.
@@
Suppose all the predictor functions ${\mathbf{h}}=(h_{1},\ldots,$ $h_{p})^{T}$ are continuous with respect to all continuous factors of $\mathbf{x}\in{\cal X}$, and $\boldsymbol{\Theta}$ is bounded. Since the number of level combinations of discrete factors is finite, there exist $M_{x}>0$ and $M_{\eta}>0$, such that, $\|{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}\|\leq M_{x}$ for all ${\mathbf{x}}\in{\cal X}$, and $|{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}|\leq M_{\eta}$ for all ${\mathbf{x}}\in{\cal X}$ and $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Since $\nu$ is continuous for commonly used GLMs (see, for example, Table 5 in the Supplementary Material of Huang et al. (2025)), there is an $M_{F}>0$, such that, $|{\mathbf{F}}_{st}({\mathbf{x}},\boldsymbol{\theta})|\leq M_{F}$ for all $s,t=1,\ldots,p$, $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and ${\mathbf{x}}\in{\cal X}$. As a direct conclusion, the Fisher information of a commonly used GLM satisfies Assumption (A3). The rest statements are direct conclusions of Theorem 2 and Corollary 1. ∎
Proof of Theorem 5.
@@
According to Theorem 2, design $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})\leq p$. By applying Theorem 2 and its proof (see their Lemma 1) of Huang et al. (2024), we obtain